# The Hidden Link Between RLHF and Contrastive Learning

### v1

@@ C.1 DV / MINE Sampling Scheme for RLHF
Positive (joint) draw.
For each prompt $x\sim D$ we sample a chosen response $y_{w}\sim\pi_{\text{c}}(\cdot\mid x)$, defining the joint measure $dP_{\pi_{\theta},\pi_{\text{c}}}=D(x)\,\pi_{\text{c}}(y\mid x)$.
Negative (product) draw.
Conditioned on the same $x$ we draw a response $y\sim\bar{\pi}(\cdot\mid x)$ from the mixed policy $\bar{\pi}=\tfrac{1}{2}\pi_{\text{c}}+\tfrac{1}{2}\pi_{\text{r}},$ yielding the product measure $dP_{\pi_{\theta}}\,dP_{\bar{\pi}}$. Since $D(x)$ cancels in any density ratio, $\pi_{\theta}$ appears only as a marginal over the prompt.
@@
C.2 Optimal Critic for the Mixed-product Bound
The optimum is a log-density ratio.
Equation (C.5) is a Fenchel–Legendre dual:
$\sup_{T}\bigl{[}\mathbb{E}_{P_{\text{joint}}}[T]-\log\mathbb{E}_{P_{\text{prod}}}[e^{T}]\bigr{]}\;=\;\mathrm{KL}\!\bigl{(}P_{\text{joint}}\,\|\,P_{\text{prod}}\bigr{)},$
whose supremum is attained when the Fenchel–Young equality holds, i.e.
$T^{\star}(x,y)\;=\;\log\frac{dP_{\text{joint}}}{dP_{\text{prod}}}(x,y).$
Substituting $P_{\text{joint}}\!=\!P_{\pi_{\theta},\pi_{\text{c}}}$ and $P_{\text{prod}}\!=\!P_{\pi_{\theta}}P_{\bar{\pi}}$ gives
$\boxed{T^{\star}(x,y)=\log\frac{\pi_{\text{c}}(y\mid x)}{\bar{\pi}(y\mid x)}}.$ (C.2)
Intuitively, $T^{\star}$ rewards samples that are more likely under $\pi_{\text{c}}$ than under the negative pool $\bar{\pi}$.
C.3 Why $T^{\star}$ Is Impractical
Evaluating $\bar{\pi}(y\mid x)$ requires two forward passes ($\pi_{\text{c}}$ and $\pi_{\text{r}}$) for every negative candidate, multiplying GPU cost. Moreover, $e^{T^{\star}}=\pi_{\text{c}}/\bar{\pi}$ can be extremely heavy-tailed when $\pi_{\text{r}}$ diverges from $\pi_{\text{c}}$, rendering the Monte-Carlo estimate of $\mathbb{E}[e^{T^{\star}}]$ numerically unstable. For these reasons optimising (C.2) is unattractive in practice.
C.4 A Restricted Critic Family
@@
$\boxed{\mathcal{F}_{\text{sub}}=\bigl{\{}T_{\theta,c}(x,y)=\log\tfrac{\pi_{\theta}(y\mid x)}{\pi_{\text{c}}(y\mid x)}+c\;\bigm{|}\;c\in\mathbb{R}\bigr{\}},}$ (C.3)
@@
Plugging $T_{\theta,c}$ into (C.5) yields the surrogate bound
$\hat{I}(\theta):=\sup_{c\in\mathbb{R}}\Bigl{[}\mathbb{E}_{P_{\pi_{\theta},\pi_{\text{c}}}}[T_{\theta,c}]-\log\mathbb{E}_{P_{\pi_{\theta}}P_{\bar{\pi}}}[e^{T_{\theta,c}}]\Bigr{]}-\log 2,$ (C.4)
@@
The first inequality reflects the restriction $\mathcal{F}_{\text{sub}}\!\subset\!\mathcal{F}_{\text{all}}$; the second arises from replacing $\pi_{\text{c}}$ by $\bar{\pi}$ in the denominator. Despite this looseness, Proposition B.1 in the main text shows that the positive and negative gradients are always anti-parallel, guaranteeing a contrastive signal at every optimisation step.
Appendix D Upper Bounds on the Jensen Gap
@@ Appendix E Supplementary: Preliminaries
$\max_{\pi_{\theta}}\mathbb{E}_{\mathbf{y}\sim\pi_{\theta}(\mathbf{y}|\mathbf{x})}\left[r(\mathbf{x},\mathbf{y})\right]-\beta\mathrm{D}_{\mathrm{KL}}\left[\pi_{\theta}(\mathbf{y}|\mathbf{x})\|\pi_{\mathrm{ref}}(\mathbf{y}|\mathbf{x})\right],$ (C.6)
where $\beta>0$ is an appropriate KL penalty coefficient. RLHF typically optimizes the above objective in Equation C.6 using RL algorithms, such as PPO [49]. Although RLHF with PPO has achieved remarkable success, the training process of PPO is unstable because of the high variance of the estimates of the policy gradients [50].
@@
$\mathcal{L}_{\mathrm{RM}}(\phi;\mathcal{D})=\mathbb{E}_{(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})\sim\mathcal{D}}\left[-\log\sigma\left(r_{\phi}(\mathbf{x},\mathbf{y}_{w})-r_{\phi}(\mathbf{x},\mathbf{y}_{l})\right)\right].$ (C.7)
@@
$I(Y;Z)=\sup_{T\in\mathcal{F}}\Bigl{(}\mathbb{E}_{P_{YZ}}[T]-\log\mathbb{E}_{P_{Y}P_{Z}}[e^{T}]\Bigr{)},$ (C.8)
where $\mathcal{F}$ is any function class with finite log-moment [21, 51]. Belghazi et al. [22] propose parametrising $T_{\phi}$ by a neural network and maximising the RHS of (C.8); the resulting estimator is known as MINE.
@@
Energy-based models (EBMs) [24] define the distribution through an energy function. For $\mathbf{y}\in\mathbb{R}^{D}$, its probability density can be expressed as follows:
$p_{\theta}(\mathbf{y})=\exp(-E_{\theta}(\mathbf{y}))/Z_{\theta}(\mathbf{y}),$ (C.9)
where $E_{\theta}(\mathbf{y}):\mathbb{R}^{D}\rightarrow\mathbb{R}$ is the energy function, mapping the data point $\mathbf{y}$ to a scalar, and $Z_{\theta}(\mathbf{y})=\sum_{\mathbf{y}}\exp(-E_{\theta}(\mathbf{y}))$ is the unknown normalization constant [52].
Appendix F Supplementary: Experimental Details
@@ F.1 Training Details
We follow the general training configurations established in SimPO. During the supervised fine-tuning (SFT) stage, we use a learning rate of $2\times 10^{-5}$. For both SFT and preference optimization, we set the batch size to 128 and the maximum sequence length to 2048. A cosine learning rate scheduler is applied with a 10% warm-up ratio over a single epoch. We use the Adam optimizer [53] in all experiments.
@@
All training experiments are conducted using six NVIDIA A6000 GPUs (46GB each), with a total batch size of 384. We build on the publicly available alignment-handbook codebase.
F.2 The Details of Datasets
UltraFeedbackBinarized [54, 55] is a dataset comprising 64k prompts, each paired with four completions generated by various open-source and proprietary language models. GPT-4 assigns scores to these completions based on criteria such as helpfulness, honesty, and other qualitative metrics. To form binary preference pairs, the completion with the highest average score is selected as the preferred response, while one of the remaining three completions is randomly chosen as the rejected response.
F.3 Evaluation Tasks: Math and Reasoning Benchmarks
@@
HendrycksMath [56] consists of 12,500 high school mathematics problems spanning algebra, geometry, and number theory. Each problem is accompanied by a step-by-step solution, targeting rigorous symbolic reasoning.
MinervaMath [57] evaluates quantitative reasoning over STEM domains using university-level mathematics, science, and engineering problems. It emphasizes complex problem solving in structured formats.
MultiMedQA [58] is a multi-task benchmark for medical question answering, aggregating datasets from professional exams, clinical research, and consumer health queries. It measures both factual correctness and clinical alignment.
MathQA [59] contains over 37k multiple-choice math word problems paired with symbolic execution programs. The dataset encourages interpretable problem solving using operation-based formalisms.
GSM8K [60] includes 8,500 elementary school math word problems requiring multi-step reasoning. It is widely used to benchmark arithmetic and logical problem-solving capabilities of language models. For this task, we report results using the exact match metric under the flexible-extract setting, which allows for variations in answer formatting.
AQuA-RAT [61] focuses on algebraic word problems from the AQuA-RAT dataset, requiring rationales and supporting steps. It is part of the broader AGIEval benchmark designed to test cognitive ability in foundation models. For this task, we use normalized accuracy (acc_norm) as the evaluation metric to account for formatting variability in symbolic responses.

### v2

@@ C.1 DV / MINE Sampling Scheme for RLHF
C.2 Let’s Constrain T To A Restricted Critic Family
@@
$\boxed{\mathcal{F}_{\text{sub}}=\bigl\{T(\pi_{\theta},\pi_{\text{c}})=\log\tfrac{\pi_{\theta}(y\mid x)}{\pi_{\text{c}}(y\mid x)}+c\;\bigm|\;c\in\mathbb{R}\bigr\},}$ (C.6)
@@
Plugging log-ratio $T$ into equation C.5 yields the surrogate bound
$\hat{I}(\theta):=\sup_{\theta}\Bigl[\mathbb{E}_{P_{\pi_{\theta},\pi_{\text{c}}}}[T(\pi_{\theta},\pi_{\text{c}})]-\log\mathbb{E}_{P_{\pi_{\theta}}P_{\bar{\pi}}}[e^{T(\pi_{\theta},\bar{\pi})}]\Bigr]-\log 2,$ (C.7)
@@
The first inequality reflects the restriction $\mathcal{F}_{\text{sub}}\!\subset\!\mathcal{F}_{\text{all}}$; the second arises from replacing $\pi_{\text{c}}$ by $\bar{\pi}$ in the denominator. Despite this looseness, Appendix C shows that the positive and negative gradients are always anti-parallel, guaranteeing a contrastive signal at every optimisation step.
Appendix D The JSD estimator yields substantially lower variance than DV/MINE
Subsequent research has shown that the DV-based objective used in MINE suffers from significant instability in practice (Poole et al., 2019). Specifically, its variance can grow exponentially with the true mutual information (MI) (Song & Ermon, 2020), leading to exploding gradients and requiring very large batch sizes to maintain stable training (Guo et al., 2022). Poole et al. (2019) further demonstrate that DV-style bounds like MINE exhibit low bias but extremely high variance, causing erratic gradient updates.
In the work (McAllester & Stratos, 2020), it is emphasized that accurate MI estimation is intrinsically challenging: high bias or high variance can lead to overfitting of spurious correlations. Thus, in contrastive learning, stable MI lower bounds (e.g., MINE or the Jensen–Shannon (JS) bound) should be prioritized over maximizing numerical MI estimates.
Additionally, Sordoni et al. (2021) point out that the MINE bound is upper-bounded by $\log K$, where $K$ is the number of negative samples. When the true MI greatly exceeds $\log K$, the bound severely underestimates MI.
Collectively, these studies (Guo et al., 2022; McAllester & Stratos, 2020) highlight that in regimes with few negative samples, MINE suffers from prohibitively large variance (Poole et al., 2019), leading to unstable gradients. Therefore, replacing MINE with a variance-reduced alternative, such as the JSD estimator (Wen et al., 2020), is crucial for reliable optimization.
The left panel of Figure 4 compares the DV/MINE and JSD estimators with a single negative sample ($K=1$). While both estimators qualitatively capture the ground-truth MI trend, JSD does not consistently outperform MINE in terms of accuracy. However, when gradient stability is prioritized over precise MI estimation, JSD demonstrates a clear advantage (Shrivastava et al., 2023). To further investigate this, we analyze the gradient variance induced by each estimator, with the right panel of Figure 4 showing the gradient variance computed from the first-layer weights of a three-layer MLP trained for each estimator’s objective. Under conditions with limited negative samples, JSD exhibits significantly greater stability than MINE, making it particularly suitable for training objectives, such as Alignment, that rely on contrastive MI estimates with sparse negative sampling (Chen et al., 2024).
Appendix E Upper Bounds on the Jensen Gap
@@ Appendix F Preliminaries
$\max_{\pi_{\theta}}\mathbb{E}_{\mathbf{y}\sim\pi_{\theta}(\mathbf{y}|\mathbf{x})}\left[r(\mathbf{x},\mathbf{y})\right]-\beta\mathrm{D}_{\mathrm{KL}}\left[\pi_{\theta}(\mathbf{y}|\mathbf{x})\|\pi_{\mathrm{ref}}(\mathbf{y}|\mathbf{x})\right],$ (F.1)
where $\beta>0$ is an appropriate KL penalty coefficient. RLHF typically optimizes the above objective in Equation F.1 using RL algorithms, such as PPO (Schulman et al., 2017). Although RLHF with PPO has achieved remarkable success, the training process of PPO is unstable because of the high variance of the estimates of the policy gradients (Engstrom et al., 2020).
@@
$\mathcal{L}_{\mathrm{RM}}(\phi;\mathcal{D})=\mathbb{E}_{(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})\sim\mathcal{D}}\left[-\log\sigma\left(r_{\phi}(\mathbf{x},\mathbf{y}_{w})-r_{\phi}(\mathbf{x},\mathbf{y}_{l})\right)\right].$ (F.2)
@@
$I(Y;Z)=\sup_{T\in\mathcal{F}}\Bigl(\mathbb{E}_{P_{YZ}}[T]-\log\mathbb{E}_{P_{Y}P_{Z}}[e^{T}]\Bigr),$ (F.3)
where $\mathcal{F}$ is any function class with finite log-moment (Donsker & Varadhan, 1983; Belghazi et al., 2018b). Belghazi et al. (2018a) propose parametrising $T_{\phi}$ by a neural network and maximising the RHS of equation F.3; the resulting estimator is known as MINE.
@@
Energy-based models (EBMs) (Lecun et al., 2006) define the distribution through an energy function. For $\mathbf{y}\in\mathbb{R}^{D}$, its probability density can be expressed as follows:
$p_{\theta}(\mathbf{y})=\exp(-E_{\theta}(\mathbf{y}))/Z_{\theta}(\mathbf{y}),$ (F.4)
where $E_{\theta}(\mathbf{y}):\mathbb{R}^{D}\rightarrow\mathbb{R}$ is the energy function, mapping the data point $\mathbf{y}$ to a scalar, and $Z_{\theta}(\mathbf{y})=\sum_{\mathbf{y}}\exp(-E_{\theta}(\mathbf{y}))$ is the unknown normalization constant (Song & Kingma, 2021).
Appendix G Experimental Details
@@ G.1 Training Details
CPO  $\beta\in[0.1,0.01]$ MIO  $\beta\in[5.0,7.5,8.75,10.0]$
We follow the general training configurations established in SimPO. During the supervised fine-tuning (SFT) stage, we use a learning rate of $2\times 10^{-5}$. For both SFT and preference optimization, we set the batch size to 128 and the maximum sequence length to 2048. A cosine learning rate scheduler is applied with a 10% warm-up ratio over a single epoch. We use the Adam optimizer (Kingma & Ba, 2014) in all experiments.
@@
The mistral-7b-sft experiments are conducted using six NVIDIA A6000 GPUs (46GB each), with a total batch size of 384, while the llama and qwen experiments were performed on eight NVIDIA H20 GPUs. We build on the publicly available alignment-handbook codebase.
G.2 The Details of Datasets
UltraFeedbackBinarized (Cui et al., 2023a; Tunstall et al., 2023b) is a dataset comprising 64k prompts, each paired with four completions generated by various open-source and proprietary language models. GPT-4 assigns scores to these completions based on criteria such as helpfulness, honesty, and other qualitative metrics. To form binary preference pairs, the completion with the highest average score is selected as the preferred response, while one of the remaining three completions is randomly chosen as the rejected response.
G.3 Evaluation Tasks: Math and Reasoning Benchmarks
@@
HendrycksMath (Hendrycks et al., 2021b) consists of 12,500 high school mathematics problems spanning algebra, geometry, and number theory. Each problem is accompanied by a step-by-step solution, targeting rigorous symbolic reasoning.
MinervaMath (Lewkowycz et al., 2022b) evaluates quantitative reasoning over STEM domains using university-level mathematics, science, and engineering problems. It emphasizes complex problem solving in structured formats.
MultiMedQA (Singhal et al., 2023) is a multi-task benchmark for medical question answering, aggregating datasets from professional exams, clinical research, and consumer health queries. It measures both factual correctness and clinical alignment.
MathQA (Amini et al., 2019b) contains over 37k multiple-choice math word problems paired with symbolic execution programs. The dataset encourages interpretable problem solving using operation-based formalisms.
GSM8K (Cobbe et al., 2021b) includes 8,500 elementary school math word problems requiring multi-step reasoning. It is widely used to benchmark arithmetic and logical problem-solving capabilities of language models. For this task, we report results using the exact match metric under the flexible-extract setting, which allows for variations in answer formatting.
AQuA-RAT (Cui et al., 2023b) focuses on algebraic word problems from the AQuA-RAT dataset, requiring rationales and supporting steps. It is part of the broader AGIEval benchmark designed to test cognitive ability in foundation models. For this task, we use normalized accuracy (acc_norm) as the evaluation metric to account for formatting variability in symbolic responses.